OpenStudy (anonymous):

A bus and a car have an inelastic head-on collision. The bus has a mass of 1.5 times 10^3 kilograms and an initial velocity of +20 meters/second. The car has a mass of 9.5 times 10^2 kilograms and an initial velocity of -26 meters/second. What is their total momentum after the collision?

OpenStudy (anonymous):

u know what is inelastic collision

OpenStudy (anonymous):

no

OpenStudy (anonymous):

An inelastic collision, in contrast to an elastic collision, is a collision in which kinetic energy is not conserved. In collisions of macroscopic bodies, some kinetic energy is turned into vibrational energy of the atoms, causing a heating effect, and the bodies are deformed.

OpenStudy (anonymous):

shamim is momentum conserved in inelastic collision

OpenStudy (shamim):

Ya

OpenStudy (anonymous):

then apply momentum conservation

OpenStudy (anonymous):

okay so i apply momentum conversation and that gives me my answer?

OpenStudy (anonymous):

yes u will get the final momentum,

OpenStudy (anonymous):

3.0 times 10 ^3 is this correct?

OpenStudy (shamim):

Total momentum=m1*u1+m2*u2 =1.5*10^3*20+9.5*10^2*(-26)=?